OpenStudy (anonymous):

The effective annual interest rate on an account compounded continuously is 4.21%. Estimate the nominal interest rate

OpenStudy (perl):

e^r = ( 1+.0421*1) , i believe

OpenStudy (perl):

( 1 + r/n)^(n*t) is the compound interest formula

OpenStudy (perl):

ok lets use 1 + r = ( 1 + i/n)^n, where i is the nominal rate, and r is effective rate

OpenStudy (perl):

because the 1 + r is the total balance after 1 year, r is the effective interest

OpenStudy (perl):

for a 1 dollar initial deposit

OpenStudy (perl):

and the right side is the nominal interest rate (that which is being compounded )

OpenStudy (perl):

so solve 1 + .0421 = e^r

OpenStudy (perl):

sorry i messed up again

OpenStudy (perl):

e^i = 1 + .0421

OpenStudy (perl):

so the nominal rate is 4.12%

OpenStudy (perl):

i = ln ( 1 + .0421 )

OpenStudy (perl):

i = .04123